# Deep Class-Incremental Learning: A Survey

### v1

@@ 3 Class-Incremental Learning: Taxonomy
There are numerous works addressing class-incremental learning in recent years, raising a heated discussion among machine learning and computer vision society. We organize these methods taxonomically from three aspects, i.e., data-centric, model-centric, and algorithm-centric, as shown in Table I. Data-centric methods concentrate on solving CIL with exemplars, which can be further classified into data replay and data regularization. Model-centric methods either regularize the model parameters from drifting away or expand the network structure for stronger representation ability. Lastly, algorithm-centric methods either utilize knowledge distillation to resist forgetting or rectify the bias in the CIL model. We list the representative methods chronologically in Figure 3 to show the research focus of different periods. In the following sections, we will discuss CIL methods from these three aspects.
3.1 Data-Centric Class-Incremental Learning
Facing the notorious catastrophic forgetting, data-centric methods seek help from extra data, e.g., exemplars. An intuitive way is to utilize former data for rehearsal, which enables the model to review former classes and resist forgetting. On the other hand, some works build regularization terms with the extra data, aiming to control the optimization direction to avoid catastrophic forgetting.
3.1.1 Data Replay
‘Replay’ is important in human cognition system [119] — a student facing final exams shall go over the textbooks to recall former memory and knowledge. This phenomenon can also be extended to the network training process, where a network can overcome catastrophic forgetting by revisiting former exemplars. We denote the sequential training process finetuning the model with $\mathcal{D}^{b}$ as ‘Finetune,’ whose loss function can be denoted as:
$\mathcal{L}$ $=\sum_{({\bf x},y)\in\mathcal{D}^{b}}\ell(f({\bf x}),y)\,$, (4)
where $\ell(\cdot,\cdot)$ measures the discrepancy between inputs, e.g., cross-entropy loss. Finetune is also known as the baseline method for class-incremental learning since it only concentrates on learning the new concepts in the current task. Consequently, it suffers severe forgetting since the model pays no attention to going over former ones. To this end, an intuitive way [32] is to save an extra exemplar set $\mathcal{E}$ (as defined in Definition 2) and include it into the model updating process:
@@ 3.1.1 Data Replay
Utilizing the exemplar set for rehearsal is simple yet effective, which leads to numerous following works. The exemplar sampling process is also similar to the active learning protocol [120], and some works propose corresponding sampling measures to select the informative ones. For example, [39] suggests sampling exemplars with high prediction entropy and near the decision boundary. The model will achieve higher generalization ability by replaying these ‘hard’ exemplars. [35] proposes estimating exemplars’ uncertainty via data augmentation. They choose instances with large prediction diversity by aggregating the predictions of multiple augmented instances. Similarly, [40] proposes to sample exemplars with a greedy strategy in online incremental learning. It proves that exemplar selection is equivalent to maximizing the diversity of exemplars with parameters gradient as the feature. If there are no explicit task boundaries, [41, 42] explore the reservoir sampling process to ensure the exemplars are i.i.d. sampled. [43] proposes to formulate the replay process into a bi-level optimization and keep intact predictions on some anchor points of past tasks. [44] introduces data replay in prototypical network [45], and utilizes the exemplars as pseudo-prototypes for embedding evaluation.
@@
Figure 3: The roadmap of class-incremental learning. We organize representative methods chronologically to show the concentration at different stages. Different colors of these methods denote the sub-categories in Table I. Knowledge distillation and data replay dominated the research before the year 2021, while model rectify and dynamic networks became popular after 2021.
@@
Since exemplars are raw images, directly saving a set of instances may consume enormous memory costs. To this end, several works are proposed to build a memory-efficient replay buffer [22]. [46] argues that extracted features $\phi({\bf x})$ are with lower dimension than raw images $\bf x$} and proposes to save features in the exemplar set to release the burden. Similarly, [47] proposes to keep low-fidelity images instead of raw images to save memory budget. However, since the distributions of extracted features and low-fidelity images may differ from the raw images, an extra adaptation process is needed for these methods, adding to the algorithm’s complexity.
@@
The above-mentioned methods achieve competitive performance by replaying former instances in the memory. Moreover, apart from directly saving instances in the exemplar set, generative models show the potential to model the distribution and generate instances [121, 122, 123], which have also been applied to class-incremental learning. We denote the aforementioned methods directly saving instances for replay as ‘direct replay,’ and the methods utilize generative models as ‘generative replay.’
@@
There often exist two models in generative replay-based CIL, i.e., the generative model for data generation and the classification model for prediction. GR [48] firstly proposes to utilize the generative adversarial network (GAN) [121] in CIL. In each updating process, it utilizes the GAN to generate the instances from former classes and then updates GAN and classification model with both old and new classes. ESGR [49] extends GR by saving extra exemplars. It also proposes to train the separate GAN for each incremental task, which does not require updating GAN incrementally. [50] extends GR by introducing the dynamic parameter generator for model adaptation at test time. FearNet [51] explores brain-inspired CIL and uses a dual-memory system in which new memories are consolidated from a network for recent memories. Recently, [52, 53, 54] explore the application of conditional GAN [124] in CIL, and [55, 56] adopt variational auto-encoders (VAE) [122] to model data distribution. Similarly, [57, 58] model each class into a Gaussian distribution and sample instances directly from the class center.
However, the performance of generative replay methods relies on the quality of generated data, which may fail in complex, large-scale inputs [125]. On the other hand, sequentially updating the generative model also suffers the catastrophic forgetting problem [126, 127, 128], which in turn increases the forgetting in incremental model training.
3.1.2 Data Regularization
@@
Apart from directly replaying former data, another group of works tries to regularize the model with former data and control the optimization direction. Since learning new classes will result in the catastrophic forgetting of old ones, the intuitive idea is to make sure that optimizing the model for new classes will not hurt former ones. GEM [59] aims to find the model that satisfies:
@@
where $g,g_{old}$ denotes the gradients of the current updating step and exemplar set, respectively. Eq. 7 requires the angle between gradients to be acute. If all the inequality constraints are satisfied, then the proposed parameter update is unlikely to increase the loss at previous tasks. However, if violations occur, GEM proposes to project the gradients $g$ to the closest gradient $\tilde{g}$ satisfying the constraints. GEM further transforms the optimization into a Quadratic Program (QP) problem. However, since the regularization in Eq. 7 is defined among all exemplars, it requires calculating loss among exemplar set and solving the QP problem in each optimization step. Hence, optimizing GEM is very time-consuming. To this end, A-GEM [60] is proposed to speed up the optimization by relaxing the constraints in Eq. 7 into a random batch. A similar idea is also adopted in [40].
@@
There are other methods to address the regularization problem with exemplars. For example, Adam-NSCL [61] proposes to sequentially optimize network parameters by projecting the candidate parameter update into the approximate null space of all previous tasks. OWM [62] only allows weights modification in the direction orthogonal to the subspace spanned by all previously learned inputs. LOGD [63] further decomposes the gradients into shared and task-specific ones. In model updating, the gradient should be close to the gradient of the new task, consistent with the gradients shared by all old tasks, and orthogonal to the space spanned by the gradients specific to the old tasks.
3.1.3 Discussions about Data-Centric Methods
Data-centric CIL algorithms mainly concentrate on resisting forgetting with the help of former data. Replay is a simple yet effective strategy, which has been widely applied to image-based camera localization [129], semantic segmentation [130, 131, 132], video classification [133], and action recognition [95]. However, since the exemplar set only saves a tiny portion of the training set, data replay may suffer the overfitting problem [59, 134]. Simultaneously, the data-imbalance problem [135, 136] also occurs due to the gap between the few-shot exemplars and the many-shot training set. Several works address this problem with balanced sampling [21, 112]. On the other hand, data regularization methods rely on specific assumptions, e.g., treating the loss of exemplars as the indicator of forgetting. These assumptions may not stand in some cases [60], which results in poor performance.
Lastly, both of these methods rely on saving exemplars from the history, which may violate user privacy in some cases [137]. Hence, when developing class-incremental learning in privacy-sensitive scenarios, we should consider learning in an exemplar-free [138, 57, 139, 107] manner.
Figure 4: Illustration of network structure evolving in backbone expansion. Left: DER [20] expands a new backbone per incremental task. Middle: FOSTER [21] adds an extra model compression stage, which maintains limited model storage. Right: MEMO [22] decouples the network structure and only expands specialized blocks.
3.2 Model-Centric Class-Incremental Learning
Model-centric CIL methods mainly concentrate on model evolvement in the learning process. For example, dynamic networks assume the network’s capacity is finite and propose expanding the network when needed to enhance representation ability. On the other hand, parameter regularization methods estimate the importance of parameters and regularize important ones to prevent them from drifting away.
3.2.1 Dynamic Networks
@@
Deep neural networks are proven to produce task-specific features [140, 141]. For example, when the training dataset contains ‘cars,’ the model tends to depict the wheels and windows. However, if the model is updated with new classes containing ‘cats,’ the features would be adapted for beards and stripes. Since the capacity of a model is limited, adapting to new features will result in the overwriting of old ones and forgetting [64]. Hence, utilizing the extracted features for beards and strides is inefficient for recognizing a car. To this end, dynamic networks are designed to dynamically adjust the model’s representation ability to fit the evolving data stream. There are several ways to expand representation ability, and we divide them into three sub-groups, i.e., neuron expansion, backbone expansion, and prompt expansion.
@@
Early works focus on neuron expansion, which adds neurons when the representation ability is not enough to capture new classes. DEN [64] formulates the adjusting process into selection, expansion, duplication, and elimination. Facing a new incremental task, the model first selectively retrains the neurons that are relevant to this task. If the retrained loss is still above some threshold, DEN considers expanding new neurons top-down and eliminating the useless ones with group-sparsity regularization. Afterward, it calculates the neuron-wise drift and duplicates neurons that drift too much from the original values. Apart from heuristically expanding and shrinking the network structure, RCL [65] formulates the network expansion process into a reinforcement learning problem and searches for the best neural architecture for each incoming task. Similarly, Neural Architecture Search (NAS) [142] is also adopted to find the optimal structure for each of the sequential tasks [66].
@@
Expanding neurons shows competitive results with expandable representation. Correspondingly, several works try to duplicate the backbone network for stronger representation ability. PNN [67] proposes learning a new backbone for each new task and fixing the former in incremental learning. It also adds layer-wise connections between old and new models to reuse former features. Expert Gate [68] also expands the backbone per incremental task, while it requires learning an extra gate to map the instance to the most suitable pathway during inference. To release the expansion cost, P&C [69] suggests a progression-compression protocol — it first expands the network to learn representative representations. Afterward, a compression process is conducted to control the total budget. [70, 71] maintain a dual-branch network for class-incremental learning, one for fast adaptation and one for slow adaption. Recently, DER [20] has been proposed to address the CIL problem. Similar to PNN, it expands a new backbone when facing new tasks and aggregates the features with a larger FC layer. Take the second incremental task for an example, where the model output is aggregated as:
@@
DER has achieved powerful performance with humble forgetting. However, saving a backbone per task requires numerous memory size, and many works are proposed to obtain expandable features with a limited memory budget. FOSTER [21] formulates the learning process in Eq. 9 as a feature-boosting [143] problem. It argues that not all expanded features are needed for incremental learning and need to be integrated to reduce redundancy. For example, suppose old classes contain ‘tigers,’ and new classes contain ‘zebras.’ In that case, the stripe will be a useful feature that could be extracted by both old and new backbones. Under such circumstances, forcing the new backbone to extract the same kind of features is less effective for recognition. Hence, FOSTER adds an extra model compression process by knowledge distillation [144]:
@@
Recently, MEMO [22] has been proposed to address the memory problem in CIL, aiming to enable model expansion with the least budget cost. It finds that in class-incremental learning, shallow layers of different models are similar, while deep layers are diverse. In other words, shallow layers are more generalizable, while deep layers are specific to the task, making expanding shallow layers less memory-efficient for CIL. Hence, MEMO proposes to decouple the backbone at middle layers: $\phi({\bf x})=\phi_{s}(\phi_{g}({\bf x}))$, where specialized block $\phi_{s}$ corresponds to the deep layers in the network, while generalized block $\phi_{g}$ corresponds to the rest shallow layers. Compared to DER, MEMO only expands specialized blocks $\phi_{s}$, and transforms Eq. 9 into:

### v2

@@ 3 Class-Incremental Learning: Taxonomy
There are numerous works addressing CIL in recent years, raising a heated discussion among machine learning and computer vision society. We organize these methods taxonomically from seven aspects, as shown in Table I. Specifically, data replay and data regularization concentrate on solving CIL with exemplars, either by revisiting former instances or using them as indicators to regularize model updating. Dynamic networks expand the network structure for stronger representation ability, while parameter regularization-based methods regularize the model parameters to prevent them from drifting away to resist forgetting. Moreover, knowledge distillation builds the mapping between incremental models to resist forgetting, and model rectification aims to reduce the biased prediction of incremental learners. Template-based classification aims to transform inference into query-template matching. We list the representative methods chronologically in Figure 3 to show the research focus of different periods.
Note that the classification rules of these categories are based on the ‘key point’ (or the special focus) of each algorithm. With the rapid development of class-incremental learning, some techniques are becoming the well-acknowledged baseline to be shared among a variety of algorithms. Hence, these seven categories are not mutually exclusive, and there is no strict boundary between them. We organize them into several categories to enhance a holistic understanding of CIL from a specific perspective. In the following sections, we will discuss CIL methods from these aspects.
3.1 Data Replay
‘Replay’ is important in human cognition system [125, 126, 127] — a student facing final exams shall go over the textbooks to recall former memory and knowledge. This phenomenon can also be extended to the network training process, where a network can overcome catastrophic forgetting by revisiting former exemplars. Correspondingly, an intuitive way [128, 82] is to save an extra exemplar set $\mathcal{E}$ (as defined in Definition 2) and include it into the model updating process:
@@ 3.1 Data Replay
Hot to construct the exemplar set? Utilizing the exemplar set for rehearsal is simple yet effective, which leads to numerous following works. The exemplar sampling process is also similar to the active learning protocol, and some works propose corresponding sampling measures to select the informative ones. For example, [33] suggests sampling exemplars with high prediction entropy and near the decision boundary. The model will achieve higher generalization ability by replaying these ‘hard’ exemplars. [26] proposes estimating exemplars’ uncertainty via data augmentation. They choose instances with large prediction diversity by aggregating the predictions of multiple augmented instances. Similarly, [34] proposes to sample exemplars with a greedy strategy in online incremental learning. It proves that exemplar selection is equivalent to maximizing the diversity of exemplars with parameters gradient as the feature. Without explicit task boundaries, [35] explores the reservoir sampling process to ensure the exemplars are i.i.d. sampled. [36] formulates the replay process into a bi-level optimization and keeps intact predictions on some anchor points of past tasks. [37] introduces data replay in prototypical network [129], and utilizes the exemplars as pseudo-prototypes for embedding evaluation. Mnemonics [40] proposes a way to parameterize exemplars and optimize them in a meta-learning manner. The framework is trained through bi-level optimizations, i.e., model-level and exemplar-level, which can be combined with various CIL algorithms.
[…]
@@
Figure 3: The roadmap of class-incremental learning. We organize representative methods chronologically to show the concentration at different stages. Different colors of these methods denote the sub-categories in Table I. Knowledge distillation and data replay dominated the research before 2021, while model rectify and dynamic networks became popular after 2021.
@@
Memory-efficient memory: Since exemplars are raw images, directly saving a set of instances may consume enormous memory costs. To this end, several works are proposed to build a memory-efficient replay buffer [17]. [38] argues that extracted features have lower dimensions than raw images and proposes saving features in the exemplar set to release the burden. Similarly, [39] proposes keeping low-fidelity images instead of raw ones. However, since the distributions of extracted features and low-fidelity images may differ from the raw images, an extra adaptation process is needed for these methods, adding to the algorithm’s complexity.
@@
Generative replay: The above-mentioned methods achieve competitive performance by replaying former instances in the memory. Apart from directly saving instances in the exemplar set, generative models show the potential to model the distribution and generate instances [153, 154], which have also been applied to class-incremental learning. We denote the aforementioned methods directly saving instances for replay as ‘direct replay,’ and the methods utilizing generative models as ‘generative replay.’
@@
There often exist two models in generative replay-based CIL, i.e., the generative model for data generation and the classification model for prediction. GR [41] firstly proposes to utilize the generative adversarial network (GAN) [153] in CIL. In each updating process, it utilizes the GAN to generate the instances from former classes and then updates GAN and classification model with both old and new classes. ESGR [42] extends GR by saving extra exemplars. It also proposes to train a separate GAN for each incremental task, which does not require updating GAN incrementally. [43] extends GR by introducing the dynamic parameter generator for model adaptation at test time. FearNet [44] explores brain-inspired CIL and uses a dual-memory system in which new memories are consolidated from a network for recent memories. Recently, [45, 46, 47] explore the application of conditional GAN [155] in CIL, and [48] adopts variational auto-encoders (VAE) [154] to model data distribution. Similarly, [49, 50] model each class into a Gaussian distribution and sample instances directly from the class center. With the prosperity of diffusion models [156], recent works also consider using diffusion models as the data generator for high-quality samples. Correspondingly, SDDR [51] studies the use of a pre-trained diffusion model as a complementary source of data. Similarly, DDGR [52] adopts a diffusion model as the generator and calculates an instruction operator through the classifier to instruct the generation of samples. There are also works [157, 158] on solving the catastrophic forgetting of these diffusion models.
Discussions: Direct replay is a simple yet effective strategy, which has been widely applied to camera localization [159], semantic segmentation [160], video classification [161], and action recognition [91]. Since it directly optimizes the loss over old exemplars, it is found to help continual learners stay in the low-loss region of prior tasks during the optimization strategy [162]. However, since the exemplar set only saves a tiny portion of the training set, data replay may suffer the overfitting problem and weaken the generalizability [53, 162]. Considering repeated optimization on a small pool of data inevitably leads to tight and unstable decision boundaries, several works tackle this problem by constraining the model’s layer-wise Lipschitz constants with regard to exemplars [163] or enlarging representational variations to alleviate representation collapse [164]. Besides, the data-imbalance problem [165] also occurs due to the gap between the few-shot exemplars and the many-shot training set. Iteratively optimizing the imbalanced training set introduces extra bias in the classifier, and several works address this problem with balanced sampling [16, 110]. Finally, since direct replay requires saving exemplars of former classes, it will face privacy issues when the raw data contains face images or resource deficiency when the raw data contains images of rare animals [27]. Under such circumstances, algorithms should be designed to learn without exemplars [49] or using privacy-friendly strategies like feature replay [50].
On the other hand, the performance of generative replay methods relies on the quality of generated data. They are found to work well on simple datasets [166, 167] while failing in complex, large-scale inputs [168]. To tackle this problem, some works find generating features is much easier than generating raw images in terms of computational complexity and semantic information. Hence, they either utilize conditional GAN to generate features [169] or VAE to model the internal representations [166]. Additionally, recent advances in diffusion models reveal a promising way to generate instances with pre-trained diffusion models [51], while utilizing pre-trained models leads to an unfair comparison to other methods without extra information. Furthermore, when sequentially updating the generative model, the catastrophic forgetting phenomena can also be observed on these generative models [157, 158]. As a result, algorithms should be designed to address catastrophic forgetting in two aspects (i.e., classifier aspect and generative model aspect) when using generative replay.
3.2 Data Regularization
@@
Apart from directly replaying former data, another group of works tries to regularize the model with former data and control the optimization direction. Since learning new classes will result in the catastrophic forgetting of old ones, the intuitive idea is to ensure that optimizing the model for new classes will not hurt former ones. GEM [53] aims to find the model that satisfies:
@@
where $g,g_{old}$ denotes the gradients of the current updating step and exemplar set, respectively. Eq. 7 requires the angle between gradients to be acute. If all the inequality constraints are satisfied, then the proposed parameter update is unlikely to increase the loss of previous tasks. However, if violations occur, GEM proposes to project the gradients $g$ to the closest gradient $\tilde{g}$ satisfying the constraints. GEM further transforms the optimization into a Quadratic Program (QP) problem. However, since the regularization in Eq. 7 is defined among all exemplars, it requires calculating loss among exemplar sets and solving the QP problem in each optimization step. Hence, optimizing GEM is very time-consuming. To this end, A-GEM [54] is proposed to speed up the optimization by relaxing the constraints in Eq. 7 into a random batch. A similar idea is also adopted in [34].
@@
There are other methods to address the regularization problem with exemplars. For example, Adam-NSCL [55] proposes to sequentially optimize network parameters by projecting the candidate parameter update into the approximate null space of all previous tasks. OWM [56] only allows weight modification in the direction orthogonal to the subspace spanned by all previously learned inputs. LOGD [57] further decomposes the gradients into shared and task-specific ones. In model updating, the gradient should be close to the gradient of the new task, consistent with the gradients shared by all old tasks, and orthogonal to the space spanned by the gradients specific to the old tasks.
Discussions: Data regularization methods utilize the exemplar set in another manner, i.e., treating the loss of them as the indicator of forgetting. They assume that the loss on exemplars is consistent with the prior tasks and align the parameter updates with the direction of the exemplar set. Hence, previous knowledge can be preserved for these exemplars due to the aligned gradient direction. However, these assumptions may not stand in some cases [54], which results in poor performance. To this end, [170] gets rid of the requirement of exemplars and manually projects the gradient direction to be orthogonal with previous ones. On the other hand, some works assume the updating rule can be meta-learned [171, 172] from a series of related tasks. MER [58] regularizes the objective of data replay so that gradients on incoming examples are more likely to have transferability and less likely to have interference with respect to past examples. iTAML [173] separates the generic feature extraction module from the task-specific classifier, thereby minimizing interference and promoting a shared feature space among tasks. [174] extends data replay with adversarial attack and meta-learns an adaptive fusion module to help allocate capacity to the knowledge of different difficulties.
Moreover, since data regularization and data replay both need to save previous data in the memory, similar problems will also occur in data regularization-based methods, e.g., overfitting, generalization issues [53, 162], and privacy concerns [149]. Consequently, it would be interesting to design privacy-friendly algorithms to build the regularization term with intermediate products for real-world applications.
3.3 Dynamic Networks
@@
Deep neural networks are proven to produce task-specific features [175]. For example, when the training dataset contains ‘cars,’ the model tends to depict the wheels and windows. However, if the model is updated with new classes containing ‘cats,’ the features would be adapted for beards and stripes. Since the capacity of a model is limited, adapting to new features will result in the overwriting of old ones and forgetting [59]. Hence, utilizing the extracted features for beards and strides is inefficient for recognizing a car. To this end, dynamic networks are designed to dynamically adjust the model’s representation ability to fit the evolving data stream. There are several ways to expand representation ability, and we divide them into three sub-groups, i.e., neuron expansion, backbone expansion, and prompt expansion.
3.3.1 Neuron Expansion
@@
Early works focus on neuron expansion adding neurons when the representation ability is insufficient to capture new classes. DEN [59] formulates the adjusting process into selection, expansion, duplication, and elimination. Facing a new incremental task, the model first selectively retrains the neurons that are relevant to this task. If the retrained loss is still above some threshold, DEN considers expanding new neurons top-down and eliminating the useless ones with group-sparsity regularization. Afterward, it calculates the neuron-wise drift and duplicates neurons that drift too much from the original values. Apart from heuristically expanding and shrinking the network structure, RCL [60] formulates the network expansion process into a reinforcement learning problem and searches for the best neural architecture for each incoming task. Similarly, Neural Architecture Search (NAS) [176] is also adopted to find the optimal structure for each of the sequential tasks [61].
Figure 4: Illustration of network structure evolving in backbone expansion. Left: DER expands a new backbone per incremental task. Middle: FOSTER adds an extra model compression stage, which maintains limited model storage. Right: MEMO decouples the network structure and only expands specialized blocks.
3.3.2 Backbone Expansion
@@
Expanding neurons shows competitive results with expandable representation. Correspondingly, several works try to duplicate the backbone network for stronger representation ability. PNN [62] proposes learning a new backbone for each new task and fixing the former in incremental learning. It also adds layer-wise connections between old and new models to reuse former features. Expert Gate [63] also expands the backbone per incremental task while it requires learning an extra gate to map the instance to the most suitable pathway during inference. To release the expansion cost, P&C [64] suggests a progression-compression protocol — it first expands the network to learn representative representations. Afterward, a compression process is conducted to control the total budget. AANets [66] partially expands stable and plastic blocks and aggregates their predictions to enhance model’s representation ability. [65, 67] maintain a dual-branch network for class-incremental learning, one for fast adaptation and one for slow adaption. Recently, DER [15] has been proposed to address the CIL problem. Similar to PNN, it expands a new backbone when facing new tasks and aggregates the features with a larger FC layer. Take the second incremental task for an example, where the model output is aggregated as:
@@
However, saving a backbone per task requires numerous memory size in DER, and many works are proposed to obtain expandable features with a limited memory budget. FOSTER [16] formulates the learning process in Eq. 9 as a feature-boosting [177] problem. It argues that not all expanded features are needed for incremental learning and need to be integrated to reduce redundancy. For example, suppose old classes contain ‘tigers,’ and new classes contain ‘zebras.’ In that case, the stripe will be a useful feature that both old and new backbones could extract. Under such circumstances, forcing the new backbone to extract the same features is less effective for recognition. Hence, FOSTER adds an extra model compression process by knowledge distillation [178]:
@@
MEMO [17] addresses the memory problem in CIL, aiming to enable model expansion with the least budget cost. It finds that in CIL, shallow layers of different models are similar, while deep layers are diverse. In other words, shallow layers are more generalizable, while deep layers are specific to the task, making expanding shallow layers less memory-efficient for CIL. Hence, MEMO proposes to decouple the backbone at middle layers: $\phi({\bf x})=\phi_{s}(\phi_{g}({\bf x}))$, where specialized block $\phi_{s}$ corresponds to the deep layers in the network, while generalized block $\phi_{g}$ corresponds to the rest shallow layers. Compared to DER, MEMO only expands specialized blocks $\phi_{s}$, and transforms Eq. 9 into: